# False Discovery Rate and Localizing Power

### v2

@@ 3.1 Synthetic data
To assess the realized (empirical) fdr across procedures we used synthetic data formed by 2000 tests, which could represent voxels in a group-level result image. For each, a test statistic $z$ was simulated following a normal distribution with zero mean and unit variance. Five different scenarios were considered for independent test statistics: (i) complete null, with no synthetic effects added; (ii) positive effects only, in which 25% of the tests had their $z$ statistic increased by 3; (iii) negative effects only, in which 25% of the tests had their $z$ statistic decreased by 3; (iv) both positive and negative effects, balanced, in which 25% of the tests had their $z$ statistic increased by 3, another 25% had their $z$ statistic decreased by 3; and (v) both positive and negative effects, unbalanced, in which 10% of the tests had their $z$ statistic increased by 3, another 40% had their $z$ statistic decreased by 3. Another five scenarios (vix) were considered; these were identical to the first five except that instead of independent tests, the 2000 test statistics were made non-independent, with a correlation of 0.25. In each scenario, uncorrected p-values were computed by referring to the normal cumulative distribution function. A summary of the scenarios is shown in Table 1.
@@
Simulations for these 10 scenarios were repeated (realized) 2000 times. Each was assessed with both bh and bky. For this simulated data evaluation, we use $q=0.05$. Four different strategies were considered, described below and summarized in Table 2. Hence, we have 80 sets (10 scenarios, 2 fdr procedures, and 4 strategies) of results, each based on 2000 realizations.
Table 2: Summary of the four strategies considered for running fdr.
@@
  • Split-tails, in which the tests are separated into two disjoint subsets, one with positive ($z>0$) and another with negative ($z<0$) statistics; then the two-tailed p-values within each subset are subjected to adjustment. For the cases in which the test statistic is zero ($z=0$), the uncorrected two-tailed p-value is 1, and so is the adjusted p-value; such pin risk is extremely unlikely with continuous test statistics. After the two disjoint sets are joined, this procedure produces one complete map of adjusted p-values; alternatively, it produces two thresholds, one for positive test statistics, one for negative test statistics.
@@
The power of each of the fdr methods and the four different strategies is the ratio between the number of true discoveries by the number of tests in which true signal was added, computed for reach realization. For directional inferences (positive or negative), these refer to the results in the positive and negative sides of the statistical map; for non-directional (both sides), the directions of the test statistic and of the true signal are ignored; as described in Table 3. Thus, power here represents a distinct quantity in relation to the usual statistical power. Here it represents the sensitivity of a multiple testing procedure judged by its ability to recover a number of tests that contain true signal, that is, tests in which the null hypothesis is false.
@@
The overall empirical fdr (observed false discovery proportion) and power were computed by averaging across realizations. Confidence intervals (95%) were calculated with the Wald method (Brown et al., 2001).
@@ 3.2 Real data
To assess these four different strategies for two-tailed fdr with real data, as well as to provide an empirical comparison between bh and bky, we use a statistical map originally produced as part of the Neuroimaging Analysis Replication and Prediction Study (narpsBotvinik-Nezer et al., 2020). The narps was an effort to assess the variability of results among different analysis teams analyzing the same imaging dataset. In the study, the teams were provided with the same raw functional magnetic resonance imaging (fmri) data, collected during a gambling task. Each team analyzed the data to investigate predefined hypotheses about brain activation related to gains and losses, and employed diverse analysis workflows, with no two teams choosing identical pipelines. We use (for no particular reason) the maps produced by team 9U7M for narps Hypothesis 1. Note that the specifics of the hypothesis are not relevant to the present analysis; consult Botvinik-Nezer et al. (2020) for details about data acquisition, the predefined hypotheses, the overall narps methodology, and its data provenance. For this real data evaluation, we use $q=0.05$.
@@ 4.1 Synthetic data
Table 4: (page 4.1) Comparison of the false discovery proportions (%) among different fdr methods and correction strategies, when false positives are counted considering one side (positive or negative) or both sides of the statistical map. These results are based on synthetic data. The nominal proportion of false discoveries is $q\cdot 100\%=5\%$; their corresponding confidence intervals (95%) are shown between parentheses.
Table 5: (page 4.1) Comparison of power (%) among different fdr methods and correction strategies, when true positives are counted considering one side (positive or negative) or both sides of the statistical map. These results are based on synthetic data. The nominal proportion of false discoveries is $q\cdot 100\%=5\%$; their corresponding confidence intervals (95%) are shown between parentheses.
@@
Scenario Canonical Combined Two-tailed Split-tails Canonical Combined Two-tailed Split-tails
@@
Scenario Canonical Combined Two-tailed Split-tails Canonical Combined Two-tailed Split-tails
@@ 4.1.1 Correction strategies
In the absence of signal, that is, under the complete null hypothesis, and with independent test statistics (scenario i), the canonical method worked well when false positives were counted among observed positive test statistics only, or among observed negative test statistics only. In these cases, the empirical false discovery proportion was within the confidence intervals, being neither conservative nor anti-conservative. When false positives were counted among all test statistics (both sides, positive and negative), the observed error rate was twice as large as for each side separately, and thus, twice as large as the nominal test level. Similar behavior was observed for the split-tails method. Combined and two-tailed methods controlled the error rate at the nominal level, albeit became conservative when only one or the other direction was considered.
With only positive or only negative true signal (scenarios ii and iii), all four correction strategies produced error rates that were within the expected level for test statistics with the same sign as the signal, even if leaning heavily towards conservativeness. However, for test statistics with the opposite sign as the signal, the observed proportion of false discoveries was not controlled by a large margin, approaching 100% error rate. When both sides were considered to count false positives (that is, all tests considered), the error rate was controlled within the nominal level, leaning again towards conservativeness, albeit less heavily.
With both positive and negative true signal in a balanced scenario (scenario iv), the error rate was controlled regardless of how the error rate was counted (positive test statistics only, negative only, or both). That control, however, leaned towards conservativeness. With unbalanced amounts of signal (scenario v), which more typically represents real experiments, the error rate was not controlled in the side opposite to the one with a preponderance of true signal, with the nominal error rate lying outside the confidence interval, and conservative on the opposite side.
@@
While the same patterns above were observed for both bh and bky methods, bky had the advantage of leading to observed false discovery rates that were closer to the nominal level in the presence of signal with the same direction as the observed test statistic, as opposed to leaning towards conservatism as bh. In other words, whenever there was signal, bky was more powerful than bh. However, this also meant that the problems with excess false discoveries observed in the side opposite from where there was more signal was also exacerbated.
@@ 4.2 Real data
Figure 3: Thresholded statistical maps using different fdr methods and correction strategies for directional inference, as well as uncorrected, for the narps data. Significant results in the positive side are those greater than the threshold $t_{\text{pos}}$; in the negative side, those smaller (more negative) than $t_{\text{neg}}$; if nothing is declared significant, the threshold is set, for display purposes, at a value more extreme than the test statistic at that side, here generically represented as $\infty$. The canonical and split-tails strategies produce generally similar results; likewise, combined and two-tailed methods produce broadly similar maps. Inferences using combined or two-tailed maps should not make directional statements about the results; these are only possible in the canonical and split-tails methods. The figure also shows slight but consistent gains in power with the bky procedure when compared to the usual bh procedure. Only visualization software that allow asymmetrical thresholds can be used to display fdr-corrected results that enable directional inferences.
The uncorrected map is thresholded, at the positive side, at $t_{\text{pos}}=1-t_{\text{cdf}}^{-1}(\alpha/2,\nu)=+1.982$, where $t_{\text{cdf}}^{-1}$ is the inverse cumulative distribution function of the Student’s $t$-distribution, $\nu=107$ are the degrees of freedom for this analysis, and $\alpha=0.05$ is the test level; it is thresholded similarly at the negative side at $t_{\text{neg}}=t_{\text{cdf}}^{-1}(\alpha/2,\nu)=-1.982$. The uncorrected thresholded maps suggests the presence of signal in both positive and negative directions, with a preponderance of effects in the negative side, involving more brain regions, particularly around the cortex.
Similar computations were done for each of the correction methods in Figure 3, which shows the thresholds. The canonical strategy produced no significant results in the positive side of the map; the same for the split-tails strategy. The combined strategy resulted in tests considered significant with both positive and negative test statistics. Since all p-values (for tests investigating positive and negative effects) are pooled in the correction, the thresholds are symmetrical. Here, fewer and smaller regions are identified as significant on the negative side (where there is a preponderance of signal) and some results on the positive side that were not seen with the canonical strategy appear significant. The two-tailed strategy produced broadly similar results as with the combined method. Finally, the split-tails strategy produced results that were visually generally similar as the canonical fdr. While the differences were minute in relation to bh, bky consistently produced slightly lower thresholds, leading to a slight increase in power.
@@ 5.1 Two-tailed tests
Two-tailed tests trace their roots to the first decades of statistics as a scientific discipline. Fisher (1925, p. 47) noted that “in some cases we wish to know the probability that the deviation, known to be positive, shall exceed an observed value, whereas in other cases the probability required is that a deviation, which is equally frequently positive and negative, shall exceed an observed value”. This allows assessment of extreme outcomes in any direction, not restricting hypotheses to a specific one. Similarly, in brain imaging, there are cases in which interest lies on only one of the tails, whereas in others, the interest lies on both tails of the distribution of test statistics. Consider, for example, an fmri study where the same information is presented via visual or auditory cues: some brain areas may be selectively modulated by vision and others by hearing; interesting comparisons involve both directions, i.e., higher involvement of some regions during visual than auditory, and vice-versa. Alternatively, consider a voxel-based morphometry study that compares gray matter volume between professional singers and professional pianists: again one might expect brain regions show differences in one group while other regions show the reverse pattern.
In effect, using two-tailed tests as default has at times been recommended for neuroimaging studies: Chen et al. (2019) argued that “it is difficult to imagine a scenario where this researcher would test all voxels simultaneously for just one directional change”. However, the researcher must be aware that the amount of errors is not guaranteed within a given direction if fdr is applied across all tests. Merely applying fdr methods as integrated into neuroimaging packages such as afni, fsl, FreeSurfer, or spm — each referenced in thousands of peer-reviewed publications annually (Poldrack et al., 2019) — and then proceeding to directional inferences, does not control the error rate within each direction, as illustrated by the synthetic and real data.
@@ 5.2 Control of the error rate
The analyses using synthetic data indicate that, while fdr-based adjustment (with either bh or bky) controls the error rate in the set of tests to which the procedure is applied, should the researcher decide to make inferences about direction of effects, extra care is necessary, otherwise such inferences can be affected by excess error rate, which can well amount to up to 100% of the tests about which inference is made if true signal is present in only one side. Simple strategies such as using a two-tailed test followed by fdr, or combining the two sets of results with the directional hypotheses into one single fdr correction, do not guarantee control over the error rate for either direction, even though these control the error rate globally, i.e., when all tests are considered. Such global test, however, provides no information on the direction of effects.
@@
Alternative approaches were considered here. For example, running fdr separately for the two set of p-values with each of the directional hypotheses (the canonical strategy), or running fdr separately for the positive and negative test statistics (the split-tails strategy). The analyses demonstrate that these approaches are able to control the error rate in each direction. However, they come with an important penalty: should the null hypothesis be true everywhere, the error rate, which should be at the level $q$ under the complete null across both sides, is doubled, approaching $2q$ when both sides are considered. Should the researcher be interested in retaining the error controlled even in the complete null, then the test level should be halved. It should be noted that trying to control fdr for all subsets simultaneously would transform it into an fwer-controlling procedure (Finner and Roters, 2001).
While the simulations using synthetic data allow quantification of the observed proportion of false discoveries and how it diverges from the nominal, expected rate, real data allows for the visualization of the different patterns that may be observed in typical studies, depending on the fdr method and correction strategy adopted. With real data, the canonical and split-tails lead to broadly similar statistically significant spatial maps; likewise, the combined and two-tailed strategies were broadly similar one to another. Two-tailed tests can be seen as a particular case of multiple testing correction, since they consider both directions of an effect. Hence, it is not surprising that the the two-tailed strategy produces nearly identical results as the combined strategy, which is similar to correction across contrasts (Alberton et al., 2020).
It should be understood, however, that areas of positive signal that appeared as significant with the two-tailed and combined strategies in the narps dataset example may have emerged simply because there is a preponderance of signal in the negative side of the map; evidence for that comes from the results using synthetic data. This is because procedures that control fdr are adaptive, that is, they start with a stringent control for the smallest p-value, and proceed in steps, each contingent on the findings of the previous one, progressively loosening stringency and continuing as long as each subsequent test continue to be declared significant (Benjamini and Hochberg, 2000; Nichols, 2007). With two-tailed and combined strategies, the large amount of tests with strong signal in one side of the map (negative in our example) allows more opportunities for tests in the opposite side (positive) to be declared significant. The canonical and split-tails strategies eliminate the opportunity for eventual strong or widespread signals present in one side impact the other, as fdr is applied separately to each side.

### v3

@@ 3.1 Synthetic data
To assess the realized (empirical) fdr across procedures we used synthetic data formed by 2000 tests, which could represent voxels in a group-level result image, or any other large set of tests. For each test, a statistic $z$ was simulated following a normal distribution with zero mean and unit variance. Five different scenarios were considered for independent test statistics: (i) complete null, with no synthetic effects added; (ii) positive effects only, in which 25% of the tests had their $z$ statistic increased by 3; (iii) negative effects only, in which 25% of the tests had their $z$ statistic decreased by 3; (iv) both positive and negative effects, balanced, in which 25% of the tests had their $z$ statistic increased by 3, another 25% had their $z$ statistic decreased by 3; and (v) both positive and negative effects, unbalanced, in which 10% of the tests had their $z$ statistic increased by 3, another 40% had their $z$ statistic decreased by 3. Another five scenarios (vix) were considered; these were identical to the first five except that instead of independent tests, the 2000 test statistics were made non-independent, with a correlation of 0.25. In each scenario, uncorrected p-values were computed by referring to the normal cumulative distribution function. A summary of the scenarios is shown in Table 1.
@@
Simulations for these 10 scenarios were repeated (realized) 2000 times. Each was assessed with both bh and bky. Six different strategies were considered, described below and summarized in Table 2. Hence, we have 120 sets of results (10 scenarios, 2 fdr procedures, and 6 strategies), each based on the 2000 realizations of random test statistics. For this simulated evaluation, we use $q=0.05$.
  • Split-tails, in which the tests are separated into two disjoint sets, one with positive ($z>0$) and another with negative ($z<0$) statistics; then the two-tailed p-values within each set are subjected to adjustment. For the cases in which the test statistic is zero ($z=0$), the uncorrected two-tailed p-value is 1, and so is the adjusted p-value; such pin risk is extremely unlikely with continuous test statistics. This procedure produces one complete map of adjusted p-values after the two sets with adjusted p-valuer are joined; alternatively, it produces two thresholds, one for positive test statistics, one for negative test statistics.
  • Canonical + bb, in which adjustment is applied to one-tailed p-values in one direction, and to one-tailed p-values in the opposite direction, similarly to the canonical strategy, but using the bb procedure, that is, one set is formed by $p$ and the other by 1p1-p. This procedure produces two complete maps of adjusted p-values, one for each direction; alternatively, it produces two thresholds, one for each direction.
  • Split-tails + bb, in which the tests are separated into two disjoint sets, one with positive ($z>0$) and another with negative ($z<0$) statistics, similarly as the split-tails strategy. Then the two-tailed p-values within each subset are subjected to adjustment following the bb procedure. This strategy produces one complete map of adjusted p-values after the two sets with adjusted p-values are joined; alternatively, it produces two thresholds, one for positive test statistics, one for negative test statistics. The same considerations regarding pin risk as in the Split-tails method hold.
For the two strategies that used bb, the first stage used the Simes procedure with test level 0.05. Note that canonical and split-tails can naturally be subjected to bb as these two strategies inherently involve two sets of tests; the same does not apply to combined or two-tailed.
Table 2: Summary of the six strategies considered for running fdr.
@@
Canonical + bb $\{p^{\text{one}}_{i}\}$, $\{1-p^{\text{one}}_{i}\}$ Two*, each with $V$ tests
Split-tails + bb $\{p^{\text{two}}_{i}:z_{i}>0\}$, $\{p^{\text{two}}_{i}:z_{i}<0\}$ Two*, with a total of $V$ tests
@@
* The two calls to fdr refer to one per set (family) in the second stage of the bb procedure; the first stage uses as many calls to fdr or to any other screening procedure as there are sets.
@@
The power of each of the fdr methods and the six different strategies is the ratio between the number of true discoveries by the number of tests in which true signal was added, computed for reach realization. For directional inferences (positive or negative), these refer to the results in the positive and negative sides of the statistical map; for non-directional (both sides), the directions of the test statistic and of the true signal are ignored; as described in Table 3. Thus, power here represents a distinct quantity in relation to the usual statistical power. Here it represents the sensitivity of a multiple testing procedure judged by its ability to recover a number of tests that contain true signal, that is, tests in which the null hypothesis is false.
@@
The overall empirical fdr (observed false discovery proportion) and power were computed by averaging across realizations. Confidence intervals (95%) were calculated with the Wald method (Brown2001).
@@ 3.2 Real data
To assess these six different strategies for two-tailed fdr with real data, as well as to provide an empirical comparison between bh and bky, we use a statistical map originally produced as part of the Neuroimaging Analysis Replication and Prediction Study (narpsBotvinik-Nezer2020). The narps was an effort to assess the variability of results among different analysis teams analyzing the same imaging dataset. In the study, the teams were provided with the same raw functional magnetic resonance imaging (fmri) data, collected during a gambling task. Each team analyzed the data to investigate predefined hypotheses about brain activation related to gains and losses, and employed diverse analysis workflows, with no two teams choosing identical pipelines. We use (for no particular reason) the maps produced by team 9U7M for narps Hypothesis 1. Note that the specifics of the hypothesis are not relevant to the present analysis; consult Botvinik-Nezer2020 for details about data acquisition, the predefined hypotheses, the overall narps methodology, and its data provenance. For this real data evaluation, we use $q=0.05$.
@@ 4.1 Synthetic data
Table 4: (page 4.1) Comparison of error rate (false discovery proportion, in %) among different fdr methods and correction strategies, when false positives are counted considering one side (positive or negative) or both sides of the statistical map. These results are based on synthetic data. The nominal proportion of false discoveries is $q\cdot 100\%=5\%$; their corresponding confidence intervals (95%) are shown between parentheses.
Table 5: (page 4.1) Comparison of power (%) among different fdr methods and correction strategies, when true positives are counted considering one side (positive or negative) or both sides of the statistical map. These results are based on synthetic data. The nominal proportion of false discoveries is $q\cdot 100\%=5\%$; their corresponding confidence intervals (95%) are shown between parentheses. For scenarios that did not include signal in the direction shown, there is no power to be calculated; these are marked with a “—”.
@@
Scn. Canonical Combined Two-tailed Split-tails Canonical+$b$ Split-tails+$b$ Canonical Combined Two-tailed Split-tails Canonical+$b$ Split-tails+$b$
@@
Scn. Canonical Combined Two-tailed Split-tails Canonical+$b$ Split-tails+$b$ Canonical Combined Two-tailed Split-tails Canonical+$b$ Split-tails+$b$
@@ 4.1.1 Correction strategies
In the absence of signal, that is, under the complete null hypothesis, and with independent test statistics (scenario i), the canonical and split-tails methods worked well when false positives were counted among observed positive test statistics only, or among observed negative test statistics only. In these cases, the empirical false discovery proportion was within the confidence intervals, being neither conservative nor anti-conservative. When false positives were counted among all test statistics (i.e., both sides), the observed error rate was twice as large as as the nominal test level. Combined and two-tailed methods controlled the error rate at the nominal level, albeit became conservative when only one or the other direction was considered; these two methods, however, became invalid other scenarios. The bb adjustment with canonical and split-tails controlled the error rate in all scenarios with either bh or bky.
With only positive or only negative true signal (scenarios ii and iii), all six correction strategies produced error rates that were within the expected level for test statistics with the same side as the signal, even if leaning heavily towards conservativeness. However, for test statistics with the opposite sign as the signal, the observed proportion of false discoveries was not controlled by a large margin, approaching 100% error rate, for the combined and two-tailed methods. When both sides were considered to count false positives (that is, all tests considered), the error rate was controlled within the nominal level, leaning again towards conservativeness, albeit less heavily.
With both positive and negative true signal in a balanced scenario (scenario iv), the error rate was controlled regardless of how the error rate was counted (positive test statistics only, negative only, or both). That control, however, leaned towards conservativeness. With unbalanced amounts of signal (scenario v), which more typically represents real experiments, the error rate was not controlled in the side opposite to the one with a preponderance of true signal, with the nominal error rate lying outside the confidence interval, while conservative on the same side.
@@
While the same patterns above were observed for both bh and bky methods, bky had the advantage of leading to observed false discovery rates that were closer to the nominal level in the presence of signal with the same direction as the observed test statistic, as opposed to leaning towards conservatism as bh. In other words, whenever there was signal, bky was more powerful than bh. However, this also meant that the problems with excess false discoveries observed in the side opposite from where there was more signal was also exacerbated (scenario v).
4.1.3 Selective inference
Under the complete null (scenario i), canonical + bb and split-tails + bb controlled the global (both sides) false discovery rate at the nominal level, eliminating the doubled error rate, yet yielding identical results as their non-bb counterparts.
@@ 4.2 Real data
Figure 3: (page 4.2) Thresholded statistical maps using different fdr methods and correction strategies for directional inference, as well as uncorrected, for the narps data. Significant results in the positive side are those greater than the threshold $t_{\text{pos}}$; in the negative side, those smaller (more negative) than $t_{\text{neg}}$; if nothing is declared significant, the threshold is set, for display purposes, at a value more extreme than the test statistic at that side, here generically represented as $\infty$. The canonical and split-tails strategies produce generally similar results; likewise, combined and two-tailed methods produce broadly similar maps. Inferences using combined or two-tailed maps should not make directional statements about the results; these are only possible in the canonical and split-tails methods. The figure also shows slight but consistent gains in power with the bky procedure when compared to the usual bh procedure. Only visualization software that allow asymmetrical thresholds can be used to display fdr-corrected results that enable directional inferences.
The uncorrected map is thresholded, at the positive side, at $t_{\text{pos}}=1-t_{\text{cdf}}^{-1}(\alpha/2,\nu)=+1.982$, where $t_{\text{cdf}}^{-1}$ is the inverse cumulative distribution function (cdf) of the Student’s $t$-distribution, $\nu=107$ are the degrees of freedom for this analysis, and $\alpha=0.05$ is the test level; it is thresholded similarly at the negative side at $t_{\text{neg}}=t_{\text{cdf}}^{-1}(\alpha/2,\nu)=-1.982$. The uncorrected thresholded maps suggests the presence of signal in both positive and negative directions, with a preponderance of effects in the negative side, involving more brain regions, particularly around the cortex.
Thresholds were calculated for the two two correction methods and six strategies, as shown in Figure 3, using the inverse cdf of Student’s $t$-distribution with the same $\nu=107$, but with test level determined by the threshold computed from the respective method and strategy. The canonical strategy produced no significant results in the positive side of the map; the same for the split-tails strategy. The combined strategy resulted in tests considered significant with both positive and negative test statistics. Since all p-values (for tests investigating positive and negative effects) are pooled in the correction, the thresholds are symmetrical. Here, fewer and smaller regions are identified as significant on the negative side (where there is a preponderance of signal) and some results on the positive side that were not seen with the canonical strategy appear significant. The two-tailed strategy produced broadly similar results as with the combined method. Finally, the split-tails strategy produced results that were visually generally similar as the canonical fdr. While the differences were minute in relation to bh, bky consistently produced slightly lower thresholds, leading to a slight increase in power. However, incorporating bb into the canonical and split-tails strategies resulted in thresholds remaining the same or further away from zero compared to their non-bb counterparts.
@@ 5.1 Two-tailed tests
Two-tailed tests trace their roots to the first decades of statistics as a scientific discipline. Fisher1925 noted that “in some cases we wish to know the probability that the deviation, known to be positive, shall exceed an observed value, whereas in other cases the probability required is that a deviation, which is equally frequently positive and negative, shall exceed an observed value”. This allows assessment of extreme outcomes in any direction, not restricting hypotheses to a specific one. Similarly, in brain imaging, there are cases in which interest lies within only one of the tails, whereas in others, the interest lies within both tails of the distribution of test statistics. Consider, for example, an fmri study where the same information is presented via visual or auditory cues: some brain areas may be selectively modulated by vision and others by hearing; interesting comparisons involve both directions, i.e., higher involvement of some regions during visual than auditory, and vice-versa. Alternatively, consider a voxel-based morphometry study that compares gray matter volume between professional singers and professional pianists: again one might expect brain regions show differences in one group while other regions show the reverse pattern.
In effect, using two-tailed tests as default has at times been recommended for neuroimaging studies, in line with the fact that most studies investigate any differences between groups, regardless of direction. In most cases there is not enough a priori information to restrict comparisons to only one direction, hence Chen2019 argued that, “it is difficult to imagine a scenario where [a] researcher would test all voxels simultaneously for just one directional change”. However, the researcher must be aware that the amount of errors is not guaranteed within a given direction if fdr is applied across all tests. Merely applying fdr methods as integrated into neuroimaging packages such as afni, fsl, FreeSurfer, or spm — each referenced in thousands of peer-reviewed publications annually (Poldrack2019) — and then proceeding to directional inferences, does not control the error rate within each direction, as illustrated here using synthetic and real data.
@@ 5.2 Control of the error rate
The analyses using synthetic data indicate that, while fdr-based adjustment (with either bh or bky) controls the error rate in the set of tests to which the procedure is applied, should the researcher decide to make inferences about direction of effects, extra care is necessary, otherwise such inferences can be affected by excess error rate. In cases where true signal is present in only one side, the error rate for the other side will be 100% if any voxels are found significant there. Simple strategies such as using a two-tailed test followed by fdr, or combining the two sets of results with the directional hypotheses into one single fdr correction, do not guarantee control over the error rate for either direction, even though these control the error rate globally, i.e., when all tests are considered. Such global tests, however, provide no information on the direction of effects, i.e., no power to localize the direction of effects.
@@
Alternative approaches were considered here. For example, running fdr separately for the two set of p-values with each of the directional hypotheses (the canonical strategy), or running fdr separately for the positive and negative test statistics (the split-tails strategy). The analyses demonstrate that these approaches are able to control the error rate in each direction. However, they come with an important penalty: should the null hypothesis be true everywhere, the error rate, which should be at the level $q$ under the complete null across both sides, is doubled, approaching $2q$ when both sides are considered. Without modification, should the researcher be interested in retaining the error controlled even in the complete null with these methods, then the test level should be halved, at the risk of loss of power. Trying to control fdr for all subsets simultaneously would transform it into an fwer-controlling procedure (Finner2001).
5.3 Selective inference
A more modest (but not less useful) goal, however, can be achieved: controlling the average false discovery error across multiple sets (families) of tests. By recasting the canonical and split-tails strategies as involving multiple such sets, the approach by Benjamini2014 (bb) comes to rescue: instead of blindly adapting the test level by dividing it by 2 to account for the doubled error rate under the complete null (observed in scenario i with canonical and split-tails), that adaptation is only applied to the extent that there is evidence allowing rejection of the null for a fraction of the sets. Moreover, the method accommodates not only two (for the two directions) but any number of ways in which the tests can be partitioned. Such partitioning can be based, for example, on regions of interest or parcellations of anatomical structures, as frequently used in brain imaging. And compared to their non-bb counterparts, the two-stage procedure does not sacrifice power, as demonstrated in Table 5 and Figure 3.